Sandy has $150in the bank. (starting point) Every month she deposits $100.
The completed number strip is:
Month 0 1 2 3 4 5 6 7
Dollars 150 250 350 450 550 650 750 850
How do you write the direct formula for this?

Answer by Alan3354(69443)

Amt = 150 + 100m
m = # of months after original deposit
Or
Amt = 50 + 100m if m=1 at the original deposit.